Integral sech √x tanh √x/√x dx


1
Expert's answer
2021-09-28T10:02:54-0400

Let "I=\int (\frac{sech(\sqrt{x}).tanh(\sqrt{x})}{\sqrt{x}}) dx\\"

Put "\sqrt{x}=t"

"\therefore \frac{1}{2\sqrt{x}}dx=dt\\
\Rightarrow \frac{1}{\sqrt{x}}dx=2.dt"

So, "I=2\int (sech(t).tanh(t))dt\\"

"=2\int (\frac{2}{ e^t + e^{-t}})(\frac{ e^t - e^{-t}}{ e^t + e^{-t}})dt\\
=4\int(\frac{ e^t - e^{-t}}{ (e^t + e^{-t})^2})dt\\"

Put "(e^t + e^{-t})=u"

"\therefore (e^t-e^{-t})dt=du\\"

So, "I=4\int (\frac{1}{u^2})du\\"

"=-\frac{4}{u}+c" where "c" is the constant of integration.

"=-\frac{4}{(e^t + e^{-t})}+c\\
=-\frac{4}{(e^{\sqrt{x}} + e^{-\sqrt{x}})}+c\\"